# Strong time-periodic solutions to a multilayered fluid-structure interaction problem

### v1

@@ 3. Transformation to a fixed domain
Without loss of generality, and following [41], we may assume that there is $\alpha>0$ with $\omega\times(-\alpha,0)\subset\Omega_{\mathrm{f}}$ and $\omega\times(0,\alpha)\subset\mathbb{R}^{3}\setminus\Omega_{\mathrm{f}}$. Moreover, we fix a cut-off function $\psi\in\mathrm{C}_{0}^{\infty}\left(\mathbb{R}^{3};\left[0,1\right]\right)$ satisfying
$\psi(z)=\left\{\begin{aligned} 1,&\enspace\text{for}\enspace z\in\left(-\frac{\alpha}{2},\frac{\alpha}{2}\right),\\ 0,&\enspace\text{for}\enspace z\in\mathbb{R}\setminus(-\alpha,\alpha).\end{aligned}\right.$
By extending $\eta$ from $\omega$ to $\mathbb{R}^{2}$ by zero, which leads to $\eta\in\mathrm{C}_{0}^{1}(\mathbb{R}^{2})$, we may define a diffeomorphism
@@
Choosing $\delta_{0}$ in (3.2) to be
(3.3) $\delta_{0}\coloneqq\frac{1}{2\|\psi^{\prime}\|_{\mathrm{L}^{\infty}(\mathbb{R})}},$
we can guarantee that $X_{\eta}$ is a $\mathrm{C}^{1}$-diffeomorphism from $\Omega_{\mathrm{f}}$ onto $\Omega_{\mathrm{f}}(t)=\Omega_{\mathrm{f}}(\eta(t))$.
If we consider time-dependent $\eta=\eta(t)$ that fulfills (3.2), with $\delta_{0}$ made precise in (3.3), for all $t\in[0,T]$, we may set $X(t,\cdot)\coloneqq X_{\eta(t)}$, and $X(t,\cdot)$ then is a $\mathrm{C}^{1}$-diffeomorphism from $\Omega_{\mathrm{f}}$ onto $\Omega_{\mathrm{f}}(t)$. In addition, for $t\in[0,T]$, we use $Y(t,\cdot)=X(t,\cdot)^{-1}$ to denote the inverse of $X(t,\cdot)$. Note that $X\in\mathrm{C}_{\mathrm{b}}^{0}([0,T];\mathrm{C}^{1}(\overline{\Omega_{\mathrm{f}}})^{3})$. Moreover, by elementary computations, for $t\in(0,T)$ and $y=(y_{1},y_{2},y_{3})\in\omega\times(-\nicefrac{{\alpha}}{{2}},\nicefrac{{\alpha}}{{2}})$, we get
$\det\nabla X(t,y)=1\enspace\text{and}\enspace\operatorname*{Cof}(\nabla X)(t,y)=\begin{pmatrix}1&0&-\partial_{y_{1}}\eta(t,y_{1},y_{2})\\ 0&1&-\partial_{y_{2}}\eta(t,y_{1},y_{2})\\ 0&0&1\end{pmatrix}.$
@@
$u(t,y)\coloneqq\operatorname*{Cof}\nabla X^{\top}(t,y)u^{\mathcal{F}}(t,X(t,y))\enspace\text{and}\enspace\pi(t,y)\coloneqq\tilde{\pi}(t,X(t,y)),\enspace\text{for}\enspace(t,y)\in(0,T)\times\Omega_{\mathrm{f}}.$
@@
(3.4) $\partial_{t}u&=\operatorname{div}\sigma_{\mathrm{f}}(u,\pi)+F(u,\pi,\eta)+f,\enspace\operatorname{div}u=0,&&t\in(0,T),\enspace x\in\Omega_{\mathrm{f}},\\ \partial_{tt}\eta+P_{\mathrm{m}}\Delta_{s}^{2}\eta-\partial_{t}\Delta_{s}\eta&=-P_{\mathrm{m}}\left(\left.(\sigma_{\mathrm{f}}(u,\pi)\mathrm{e}_{3})\right|_{\Gamma_{s}}\cdot\mathrm{e}_{3}\right)\\ &\qquad+P_{\mathrm{m}}\left(G\left(u,\pi,\eta\right)\right)\\ &\qquad+P_{\mathrm{m}}\left(\left.(\sigma_{\mathrm{s}}(d,\partial_{t}d)\mathrm{e}_{3})\right|_{\Gamma_{\mathrm{s}}}\cdot\mathrm{e}_{3}\right)+P_{\mathrm{m}}g,&&t\in(0,T),\enspace s\in\omega,\\ \int_{\omega}\eta(t,s)\,\mathrm{d}s&=0,&&t\in(0,T),\\ \partial_{tt}d&=\operatorname{div}\sigma_{\mathrm{s}}(d,\partial_{t}d)+h,&&t\in(0,T),\enspace x\in\Omega_{\mathrm{s}},\\ u(t,s,0)&=\partial_{t}\eta(t,s)\mathrm{e}_{3},&&t\in(0,T),\enspace s\in\omega,\\ d(t,x)&=\eta(t,s)\mathrm{e}_{3},&&t\in(0,T),\enspace x\in\Gamma_{\mathrm{s}},\\ u(t,x)&=0,&&t\in(0,T),\enspace x\in\Gamma_{0},\\ d(t,x)&=0,&&t\in(0,T),\enspace x\in\partial\Omega_{\mathrm{s}}\setminus\Gamma_{\mathrm{s}},\\ \eta=\nabla_{s}\eta\cdot n_{\omega}&=0,&&t\in(0,T),\enspace s\in\partial\omega,\\ u(0,\cdot)&=u(T,\cdot),&&\enspace\text{in}\enspace\Omega_{\mathrm{f}},\\ \eta(0,\cdot)&=\eta(T,\cdot)\enspace\text{and}\enspace\partial_{t}\eta(0,\cdot)=\partial_{t}\eta(T,\cdot),&&\enspace\text{in}\enspace\omega,\\ d(0,\cdot)&=d(T,\cdot)\enspace\text{and}\enspace\partial_{t}d(0,\cdot)=\partial_{t}d(T,\cdot),&&\enspace\text{in}\enspace\Omega_{\mathrm{s}}.\end{aligned}\right.$
@@
$u(t,y)=b(t,y)u^{\mathcal{F}}(t,X(t,y))\enspace\text{and}\enspace u^{\mathcal{F}}(t,x)=a(t,x)u(t,Y(t,x)).$
@@
$\mathbb{E}_{1}^{u}$ $\coloneqq\mathrm{W}^{1,p}(0,T;\mathrm{L}^{q}(\Omega_{\mathrm{f}})^{3})\cap\mathrm{L}^{p}(0,T;\mathrm{W}^{2,q}(\Omega_{\mathrm{f}})^{3}),\enspace\mathbb{E}_{1}^{\pi}\coloneqq\mathrm{L}^{p}(0,T;\mathrm{W}_{\mathrm{m}}^{1,q}(\Omega_{\mathrm{f}})),$
@@
$\mathbb{E}_{1}^{d_{1}}$ $\coloneqq\mathrm{W}^{1,p}(0,T;\mathrm{W}^{1,q}(\Omega_{\mathrm{s}})^{3})\enspace\text{and}\enspace\mathbb{E}_{1}^{d_{2}}\coloneqq\mathrm{W}^{1,p}(0,T;\mathrm{L}^{q}(\Omega_{\mathrm{s}})^{3})\cap\mathrm{L}^{p}(0,T;\mathrm{W}^{1,q}(\Omega_{\mathrm{s}})^{3}).$
@@
(3.7) $\tilde{\mathbb{E}}_{1}$$\coloneqq\bigl{\{}(u,\eta_{1},\eta_{2},d_{1},d_{2})\in\mathbb{E}_{1}^{u}\times\mathbb{E}_{1}^{\eta_{1}}\times\mathbb{E}_{1}^{\eta_{2}}\times\mathbb{E}_{1}^{d_{1}}\times\mathbb{E}_{1}^{d_{2}}:d_{1}+\delta d_{2}\in\mathrm{L}^{p}(0,T;\mathrm{W}^{2,q}(\Omega_{\mathrm{s}})^{3})\bigr{\}}\enspace\text{and}\enspace$
$\mathbb{E}_{1}$$\coloneqq\bigl{\{}v=(u,\pi,\eta_{1},\eta_{2},d_{1},d_{2}):(u,\eta_{1},\eta_{2},d_{1},d_{2})\in\tilde{\mathbb{E}}_{1}\enspace\text{and}\enspace\pi\in\mathbb{E}_{1}^{\pi}\bigr{\}}.$
The main result stated in terms of the reference configuration is now given as follows.
@@ 4. Analysis of the linearized problem
[…]
@@
[…]
A simple rescaling argument reveals that the Arendt-Bu theorem is also valid for general time intervals $(0,T)$. On the other hand, the spectral mapping theorem for generators of analytic semigroups, see e. g. [20, Cor. IV.3.12], and the fact that the maximal $\mathrm{L}^{p}$-regularity already implies the generation of an analytic semigroup yield that the spectral condition $1\in\rho(\mathrm{e}^{2\pi A})$ is equivalent with $0\in\rho(A)$.
@@
(4.2) $\partial_{t}u-\operatorname{div}\sigma_{\mathrm{f}}(u,\pi)&=f,\enspace\operatorname{div}u=0,&&t\in(0,T),\enspace y\in\Omega_{\mathrm{f}},\\ \partial_{t}\eta_{1}&=\eta_{2},&&t\in(0,T),\enspace s\in\omega,\\ \partial_{t}\eta_{2}+P_{\mathrm{m}}\Delta_{s}^{2}\eta_{1}-\Delta_{s}\eta_{2}&=-P_{\mathrm{m}}\left(\left.\sigma_{\mathrm{f}}(u,\pi)\right|_{\Gamma}\mathrm{e}_{3}\cdot\mathrm{e}_{3}\right)\\ &\qquad+P_{\mathrm{m}}\left(\left.(\sigma_{\mathrm{s}}(d_{1},d_{2})\mathrm{e}_{3})\right|_{\Gamma_{\mathrm{s}}}\cdot\mathrm{e}_{3}\right)+P_{\mathrm{m}}g,&&t\in(0,T),\enspace s\in\omega,\\ \int_{\omega}\eta_{1}(t,s)\,\mathrm{d}s&=0,&&t\in(0,T),\\ \partial_{t}d_{1}&=d_{2},&&t\in(0,T),\enspace y\in\Omega_{\mathrm{s}},\\ \partial_{t}d_{2}-\operatorname{div}\sigma_{\mathrm{s}}(d_{1},d_{2})&=h,&&t\in(0,T),\enspace y\in\Omega_{\mathrm{s}},\\ u&=P_{\mathrm{m}}(\eta_{2})\mathrm{e}_{3},&&t\in(0,T),\enspace y\in\partial\Omega_{\mathrm{f}}\setminus\Gamma_{0},\\ d_{1}&=P_{\mathrm{m}}(\eta_{1})\mathrm{e}_{3},&&t\in(0,T),\enspace y\in\Gamma_{\mathrm{s}},\\ u&=0,&&t\in(0,T),\enspace y\in\Gamma_{0},\\ d_{1}&=0,&&t\in(0,T),\enspace y\in\partial\Omega_{\mathrm{s}}\setminus\Gamma_{\mathrm{s}},\\ \eta_{1}=\nabla_{s}\eta_{1}\cdot n_{\omega}&=0,&&t\in(0,T),\enspace s\in\partial\omega,\\ u(0,\cdot)&=u(T,\cdot),&&\enspace\text{in}\enspace\Omega_{\mathrm{f}},\\ \eta_{1}(0,\cdot)&=\eta_{1}(T,\cdot)\enspace\text{and}\enspace\eta_{2}(0,\cdot)=\eta_{2}(T,\cdot),&&\enspace\text{in}\enspace\omega,\\ d_{1}(0,\cdot)&=d_{1}(T,\cdot)\enspace\text{and}\enspace d_{2}(0,\cdot)=d_{2}(T,\cdot),&&\enspace\text{in}\enspace\Omega_{\mathrm{s}}.\end{aligned}\right.$
We remark that the term $P_{\mathrm{m}}\left(\left.\sigma_{\mathrm{f}}(u,\pi)\right|_{\Gamma}\mathrm{e}_{3}\cdot\mathrm{e}_{3}\right)$ simplifies significantly due to the the divergence-free condition $\operatorname{div}u=0$ in $\Omega_{\mathrm{f}}$ as well as the fact that $u$ vanishes in horizontal direction on $\partial\Omega_{\mathrm{f}}\setminus\Gamma_{0}$ with regard to the condition $u=P_{\mathrm{m}}(\eta_{2})\mathrm{e}_{3}$. This leads to $\left.(\nabla u+(\nabla u)^{\top})\right|_{\Gamma}=0$. As a result, with the modified trace operator
(4.3) $\gamma_{\mathrm{m}}f\coloneqq P_{\mathrm{m}}(\left.f\right|_{\Gamma})=f(\cdot,0)-\frac{1}{|\omega|}\int_{\omega}f(s^{\prime},0)\,\mathrm{d}s^{\prime},\enspace\text{for}\enspace f\in\mathrm{W}^{r,q}(\Omega_{\mathrm{f}}),\enspace r>\nicefrac{{1}}{{q}},$
@@
(4.4) $-P_{\mathrm{m}}\left(\left.\sigma_{\mathrm{f}}(u,\pi)\right|_{\Gamma}\mathrm{e}_{3}\cdot\mathrm{e}_{3}\right)=\gamma_{\mathrm{m}}\pi.$

### v2

@@ 4. Transformation to a fixed domain
Without loss of generality, and following [48], we may assume that there is $\alpha>0$ with $\omega\times(-\alpha,0)\subset\Omega_{\mathrm{f}}$ and $\omega\times(0,\alpha)\subset\mathbb{R}^{3}\setminus\Omega_{\mathrm{f}}$. Moreover, we fix a cut-off function $\psi\in\mathrm{C}_{\mathrm{c}}^{\infty}\left(\mathbb{R}^{3};\left[0,1\right]\right)$ satisfying
$\psi(z)=\left\{\begin{aligned} 1,&\kern 5.0pt\text{for}\kern 5.0ptz\in\left(-\frac{\alpha}{2},\frac{\alpha}{2}\right),\\ 0,&\kern 5.0pt\text{for}\kern 5.0ptz\in\mathbb{R}\setminus(-\alpha,\alpha).\end{aligned}\right$.
By extending $\eta$ from $\omega$ to $\mathbb{R}^{2}$ by zero, which leads to $\eta\in\mathrm{C}_{\mathrm{c}}^{1}(\mathbb{R}^{2})$, we may define a diffeomorphism
@@
We choose $\delta_{0}$ in (4.2) as
(4.3) $\delta_{0}\coloneqq\frac{1}{2\|\psi^{\prime}\|_{\mathrm{L}^{\infty}(\mathbb{R})}}.$
Let $\eta=\eta(t)$ be a time-dependent plate displacement satisfying the clamped boundary conditions $\eta=0$ and $\nabla_{s}\eta\cdot n_{\omega}=0$ on $\partial\omega$, and assume that $\eta$ fulfills the smallness condition (4.2) with this choice of $\delta_{0}$ for all $t\in[0,T]$. If $\eta(t)\in\mathrm{C}^{1}(\overline{\omega})$ for each $t\in[0,T]$, we set $X(t,\cdot)\coloneqq X_{\eta(t)}$ and $X(t,\cdot)$ is a $\mathrm{C}^{1}$-diffeomorphism from $\Omega_{\mathrm{f}}$ onto $\Omega_{\mathrm{f}}(t)$ for all $t\in[0,T]$. Moreover, we observe that $\eta\in\mathrm{C}([0,T];\mathrm{B}_{qp}^{4-\nicefrac{{2}}{{p}}}(\omega))$. Under the assumptions of Theorem 3.2 on $p$ and $q$, it follows from Sobolev embeddings that $\eta\in\mathrm{C}^{1}(\overline{\omega})$ for all $t\in[0,T]$. We refer to the proof of Subsection 6.1 for details.
In addition, for $t\in[0,T]$, we use $Y(t,\cdot)=X(t,\cdot)^{-1}$ to denote the inverse of $X(t,\cdot)$. Note that $X\in\mathrm{C}_{\mathrm{b}}^{0}([0,T];\mathrm{C}^{1}(\overline{\Omega_{\mathrm{f}}})^{3})$. Moreover, by elementary computations, for $t\in(0,T)$ and $y=(y_{1},y_{2},y_{3})\in\omega\times(-\nicefrac{{\alpha}}{{2}},\nicefrac{{\alpha}}{{2}})$, we get
$\det\nabla X(t,y)=1\kern 5.0pt\text{and}\kern 5.0pt\operatorname*{Cof}(\nabla X)(t,y)=\begin{pmatrix}1&0&-\partial_{y_{1}}\eta(t,y_{1},y_{2})\\ 0&1&-\partial_{y_{2}}\eta(t,y_{1},y_{2})\\ 0&0&1\end{pmatrix}$.
@@
$u(t,y)\coloneqq\operatorname*{Cof}\nabla X^{\top}(t,y)u^{\mathcal{F}}(t,X(t,y))\kern 5.0pt\text{and}\kern 5.0pt\pi(t,y)\coloneqq\tilde{\pi}(t,X(t,y)),\kern 5.0pt\text{for}\kern 5.0pt(t,y)\in(0,T)\times\Omega_{\mathrm{f}}.$
@@
(4.4) $\left\{\begin{aligned} \partial_{t}u&=\operatorname{div}\sigma_{\mathrm{f}}(u,\pi)+F(u,\pi,\eta)+f,\kern 5.0pt\operatorname{div}u=0,&&\kern 5.0pt\text{in}\kern 5.0pt(0,T)\times\Omega_{\mathrm{f}},\\ \partial_{tt}\eta+P_{\mathrm{m}}\Delta_{s}^{2}\eta-\partial_{t}\Delta_{s}\eta&=-P_{\mathrm{m}}\left(\left.(\sigma_{\mathrm{f}}(u,\pi)\mathrm{e}_{3})\right|_{\Gamma_{0}}\cdot\mathrm{e}_{3}\right)\\ &\qquad+P_{\mathrm{m}}\left(G\left(u,\pi,\eta\right)\right)\\ &\qquad+P_{\mathrm{m}}\left(\left.(\sigma_{\mathrm{s}}(d,\partial_{t}d)\mathrm{e}_{3})\right|_{\Gamma_{0}}\cdot\mathrm{e}_{3}\right)+P_{\mathrm{m}}g,&&\kern 5.0pt\text{in}\kern 5.0pt(0,T)\times\omega,\\ \int_{\omega}\eta(t,s)\,\mathrm{d}s&=0,&&\kern 5.0pt\text{for}\kern 5.0ptt\in(0,T),\\ \partial_{tt}d&=\operatorname{div}\sigma_{\mathrm{s}}(d,\partial_{t}d)+h,&&\kern 5.0pt\text{in}\kern 5.0pt(0,T)\times\Omega_{\mathrm{s}},\\ u(t,s,0)&=\partial_{t}\eta(t,s)\mathrm{e}_{3},&&\kern 5.0pt\text{for}\kern 5.0ptt\in(0,T),\kern 5.0pts\in\omega,\\ d(t,s,0)&=\eta(t,s)\mathrm{e}_{3},&&\kern 5.0pt\text{for}\kern 5.0ptt\in(0,T),\kern 5.0pts\in\omega,\\ u(t,x)&=0,&&\kern 5.0pt\text{for}\kern 5.0ptt\in(0,T),\kern 5.0ptx\in\Gamma_{\mathrm{f}},\\ d(t,x)&=0,&&\kern 5.0pt\text{for}\kern 5.0ptt\in(0,T),\kern 5.0ptx\in\Gamma_{\mathrm{s}},\\ \eta=\nabla_{s}\eta\cdot n_{\omega}&=0,&&\kern 5.0pt\text{on}\kern 5.0pt(0,T)\times\partial\omega,\\ u(0,\cdot)&=u(T,\cdot),&&\kern 5.0pt\text{in}\kern 5.0pt\Omega_{\mathrm{f}},\\ \eta(0,\cdot)&=\eta(T,\cdot)\kern 5.0pt\text{and}\kern 5.0pt\partial_{t}\eta(0,\cdot)=\partial_{t}\eta(T,\cdot),&&\kern 5.0pt\text{in}\kern 5.0pt\omega,\\ d(0,\cdot)&=d(T,\cdot)\kern 5.0pt\text{and}\kern 5.0pt\partial_{t}d(0,\cdot)=\partial_{t}d(T,\cdot),&&\kern 5.0pt\text{in}\kern 5.0pt\Omega_{\mathrm{s}}.\end{aligned}\right$.
@@
$u(t,y)=b(t,y)u^{\mathcal{F}}(t,X(t,y))\kern 5.0pt\text{and}\kern 5.0ptu^{\mathcal{F}}(t,x)=a(t,x)u(t,Y(t,x)).$
@@
$\mathbb{E}_{1}^{u}$ $\coloneqq\mathrm{W}^{1,p}(0,T;\mathrm{L}^{q}(\Omega_{\mathrm{f}})^{3})\cap\mathrm{L}^{p}(0,T;\mathrm{W}^{2,q}(\Omega_{\mathrm{f}})^{3}),\kern 5.0pt\mathbb{E}_{1}^{\pi}\coloneqq\mathrm{L}^{p}(0,T;\mathrm{W}_{\mathrm{m}}^{1,q}(\Omega_{\mathrm{f}})),$
@@
$\mathbb{E}_{1}^{d_{1}}$ $\coloneqq\mathrm{W}^{1,p}(0,T;\mathrm{W}^{1,q}(\Omega_{\mathrm{s}})^{3})\kern 5.0pt\text{and}\kern 5.0pt\mathbb{E}_{1}^{d_{2}}\coloneqq\mathrm{W}^{1,p}(0,T;\mathrm{L}^{q}(\Omega_{\mathrm{s}})^{3})\cap\mathrm{L}^{p}(0,T;\mathrm{W}^{1,q}(\Omega_{\mathrm{s}})^{3}).$
@@
(4.7) $\tilde{\mathbb{E}}_{1}$ $\coloneqq\bigl\{(u,\eta_{1},\eta_{2},d_{1},d_{2})\in\mathbb{E}_{1}^{u}\times\mathbb{E}_{1}^{\eta_{1}}\times\mathbb{E}_{1}^{\eta_{2}}\times\mathbb{E}_{1}^{d_{1}}\times\mathbb{E}_{1}^{d_{2}}:d_{1}+\delta d_{2}\in\mathrm{L}^{p}(0,T;\mathrm{W}^{2,q}(\Omega_{\mathrm{s}})^{3})\bigr\}\kern 5.0pt\text{and}\kern 5.0pt$
$\mathbb{E}_{1}$ $\coloneqq\bigl\{v=(u,\pi,\eta_{1},\eta_{2},d_{1},d_{2}):(u,\eta_{1},\eta_{2},d_{1},d_{2})\in\tilde{\mathbb{E}}_{1}\kern 5.0pt\text{and}\kern 5.0pt\pi\in\mathbb{E}_{1}^{\pi}\bigr\}.$
The main result stated in terms of the reference configuration is now given as follows. Let us note that the difference with the spaces in Theorem 3.2 is that the the latter spaces concern the moving domain problem, i.e., they are defined by invoking the diffeomorphism $X$ from $\Omega_{\mathrm{f}}$ to $\Omega_{\mathrm{f}}(t)$. Moreover, note that the definition of a strong solution to (4.4) is analogous to Section 3 in the context of the problem on the moving domain.
@@ 5. Analysis of the linearized problem
(5.1) $\left\{\begin{aligned} w^{\prime}(t)-Aw(t)&=f(t),\kern 5.0pt\kern 5.0pt\text{for}\kern 5.0ptt\in(0,2\pi),\\ w(0)&=w(2\pi),\end{aligned}\right$.
@@
$\left\{\begin{aligned} w^{\prime}(t)-Aw(t)&=f(t),\kern 5.0pt\kern 5.0pt\text{for}\kern 5.0ptt\in(0,2\pi),\\ w(0)&=0.\end{aligned}\right$.
A simple rescaling argument reveals that the Arendt-Bu theorem is also valid for general time intervals $(0,T)$. On the other hand, the spectral mapping theorem for generators of analytic semigroups, see, e.g., [26, Cor. IV.3.12], and the fact that the maximal $\mathrm{L}^{p}$-regularity already implies the generation of an analytic semigroup yield that the spectral condition $1\in\rho(\mathrm{e}^{2\pi A})$ is equivalent with $0\in\rho(A)$.
@@
(5.2) $\left\{\begin{aligned} \partial_{t}u-\operatorname{div}\sigma_{\mathrm{f}}(u,\pi)&=f,\kern 5.0pt\operatorname{div}u=0,&&\kern 5.0pt\text{in}\kern 5.0pt(0,T)\times\Omega_{\mathrm{f}},\\ \partial_{t}\eta_{1}&=\eta_{2},&&\kern 5.0pt\text{in}\kern 5.0pt(0,T)\times\omega,\\ \partial_{t}\eta_{2}+P_{\mathrm{m}}\Delta_{s}^{2}\eta_{1}-\Delta_{s}\eta_{2}&=-P_{\mathrm{m}}\left(\left.\sigma_{\mathrm{f}}(u,\pi)\right|_{\Gamma_{0}}\mathrm{e}_{3}\cdot\mathrm{e}_{3}\right)\\ &\qquad+P_{\mathrm{m}}\left(\left.(\sigma_{\mathrm{s}}(d_{1},d_{2})\mathrm{e}_{3})\right|_{\Gamma_{0}}\cdot\mathrm{e}_{3}\right)+P_{\mathrm{m}}g,&&\kern 5.0pt\text{in}\kern 5.0pt(0,T)\times\omega,\\ \int_{\omega}\eta_{1}(t,s)\,\mathrm{d}s&=0,&&\kern 5.0pt\text{for}\kern 5.0ptt\in(0,T),\\ \partial_{t}d_{1}&=d_{2},&&\kern 5.0pt\text{in}\kern 5.0pt(0,T)\times\Omega_{\mathrm{s}},\\ \partial_{t}d_{2}-\operatorname{div}\sigma_{\mathrm{s}}(d_{1},d_{2})&=h,&&\kern 5.0pt\text{in}\kern 5.0pt(0,T)\times\Omega_{\mathrm{s}},\\ u&=P_{\mathrm{m}}(\eta_{2})\mathrm{e}_{3},&&\kern 5.0pt\text{on}\kern 5.0pt(0,T)\times\Gamma_{0},\\ d_{1}&=P_{\mathrm{m}}(\eta_{1})\mathrm{e}_{3},&&\kern 5.0pt\text{on}\kern 5.0pt(0,T)\times\Gamma_{0},\\ u&=0,&&\kern 5.0pt\text{on}\kern 5.0pt(0,T)\times\Gamma_{\mathrm{f}},\\ d_{1}&=0,&&\kern 5.0pt\text{on}\kern 5.0pt(0,T)\times\Gamma_{\mathrm{s}},\\ \eta_{1}=\nabla_{s}\eta_{1}\cdot n_{\omega}&=0,&&\kern 5.0pt\text{on}\kern 5.0pt(0,T)\times\partial\omega,\\ u(0,\cdot)&=u(T,\cdot),&&\kern 5.0pt\text{in}\kern 5.0pt\Omega_{\mathrm{f}},\\ \eta_{1}(0,\cdot)&=\eta_{1}(T,\cdot)\kern 5.0pt\text{and}\kern 5.0pt\eta_{2}(0,\cdot)=\eta_{2}(T,\cdot),&&\kern 5.0pt\text{in}\kern 5.0pt\omega,\\ d_{1}(0,\cdot)&=d_{1}(T,\cdot)\kern 5.0pt\text{and}\kern 5.0ptd_{2}(0,\cdot)=d_{2}(T,\cdot),&&\kern 5.0pt\text{in}\kern 5.0pt\Omega_{\mathrm{s}}.\end{aligned}\right$.
We remark that the term $P_{\mathrm{m}}\left(\left.\sigma_{\mathrm{f}}(u,\pi)\right|_{\Gamma_{0}}\mathrm{e}_{3}\cdot\mathrm{e}_{3}\right)$ simplifies significantly due to the the divergence-free condition $\operatorname{div}u=0$ in $\Omega_{\mathrm{f}}$ as well as the fact that $u$ vanishes in horizontal direction on $\Gamma_{0}$ with regard to the condition $u=P_{\mathrm{m}}(\eta_{2})\mathrm{e}_{3}$. This leads to $\left.(\nabla u+(\nabla u)^{\top})\right|_{\Gamma_{0}}=0$. As a result, with the modified trace operator
(5.3) $\gamma_{\mathrm{m}}f\coloneqq P_{\mathrm{m}}(\left.f\right|_{\Gamma_{0}})=f(\cdot,0)-\frac{1}{|\omega|}\int_{\omega}f(s^{\prime},0)\,\mathrm{d}s^{\prime},\kern 5.0pt\text{for}\kern 5.0ptf\in\mathrm{W}^{r,q}(\Omega_{\mathrm{f}}),\kern 5.0ptr>\nicefrac{{1}}{{q}},$
@@
(5.4) $P_{\mathrm{m}}\left(\left.\sigma_{\mathrm{f}}(u,\pi)\right|_{\Gamma_{0}}\mathrm{e}_{3}\cdot\mathrm{e}_{3}\right)=\gamma_{\mathrm{m}}\pi.$